# The $u$-invariant of function fields in one variable

### v1

@@ Abstract.
The $u$-invariant of a field is the largest dimension of an anisotropic quadratic torsion form over the field. In this article we obtain a bound on the $u$-invariant of function fields in one variable over a henselian valued field with arbitrary value group and with residue field of characteristic different from $2$. This generalises a theorem due to Harbater, Hartmann and Krashen and its extension due to Scheiderer. Their result covers the special case where the valuation is discrete.
Keywords: Quadratic form, henselian valued field, function field in one variable, quadratic form, local-global principle, field ordering, torsion form, general $u$-invariant
@@ 1. Introduction
The study of numerical field invariants in quadratic form theory was initiated by Irving Kaplansky in 1953. In [16], he introduces three quantities in ${}\mathbb{N}\cup\{\infty\}$ attached to an arbitrary field and suggests to study and compare them. One of them was later called the $u$-invariant, a term which goes back to Hanfried Lenz’ article [20] from 1963.
A field $K$ is called nonreal if $-1$ is a sum of squares in $K$, and otherwise $K$ is called real. For this introduction, let $K$ denote a nonreal field of characteristic different from $2$.
@@
As pointed out in [16, Theorem 3], one always has $u(K(\!(t)\!))=2u(K)$, where $K(\!(t)\!)$ denotes the field of Laurent series in $t$ over $K$. Consequently, and since $u(\mathbb{C})=1$, examples of fields of $u$-invariant $2^{r}$ for any given $r\in\mathbb{N}$ are easy to obtain. Incited by this observation, Kaplansky wonders in [16] whether $u(K)$ is always a power of $2$ or $\infty$. This question was answered in the negative in 1989 by Alexander Merkurjev. After giving in [24] a first construction of a field of $u$-invariant $6$, he showed in [25] that every positive even integer is the $u$-invariant of some field. Merkurjev’s construction of examples involves direct limits and produces fields which are not finitely generated over any proper subfield. When restricting to fields which are of finite transcendence degree over some familiar base field, no examples are known with a $u$-invariant that is not a power of $2$.
Let $n\in\mathbb{N}$ and let $F/K$ be a field extension of transcendence degree $n$. One may try to determine or at least bound the $u$-invariant of $F$ in terms of $n$ and properties of $K$. Well-known results of this sort are the following:
  • If $K$ is algebraically closed, then $u(F)\leqslant 2^{n}$, and equality holds if $F/K$ is finitely generated.
  • If $K$ is finite, then $u(F)\leqslant 2^{n+1}$, and equality holds if $F/K$ is finitely generated.
  • If $K$ is a complete discretely valued field with finite residue field, then $u(F)\leqslant 2^{n+2}$, and equality holds if $F/K$ is finitely generated.
In all three cases, the upper bounds are based on the study of systems of forms, while the lower bounds when $F/K$ is finitely generated follow from a classical argument using valuations. The first two results were established by Serge Lang in 1952 [18, Theorem 6, Cor.], whereas the last result was derived by David Leep in 2013 [19] from a result on systems of quadratic forms due to Roger Heath-Brown [14].
Note that in all the situations above, when $F/K$ is finitely generated, we have that $u(F)=2^{d}$ where $d$ is the cohomological dimension of $F$. This might suggest that $u(F)=2^{n+2}$ would hold as well when $K$ is a nonreal number field and $F/K$ a finitely generated extension of transcendence degree $n$. However, even for $n=1$ this is still an open problem. A more general open question is whether $u(F)$ can be bounded in terms of $u(K)$ and the transcendence degree of $F/K$.
@@
In particular, for any odd prime number $p$, applying 1.1 to the field of $p$-adic numbers $\mathbb{Q}_{p}$ with its natural valuation, one retrieves the result from [26] and [15] that $u(F)=8$ for every function field in one variable $F/\mathbb{Q}_{p}$.
Our main result provides a transfer principle for this common upper bound between a henselian valued field $K$ and its residue field, assuming the latter to be of characteristic different from $2$.
Our standard reference for the theory of valued fields is [13]. For a valuation $v$ on $K$, we denote by $Kv$ the residue field and by $vK$ the value group of $v$, and we call $v$ discrete if $vK\simeq\mathbb{Z}$. Given a valuation $v$ on $K$, we call the pair $(K,v)$ a valued field. A valuation $v$ on $K$ is called henselian if, for every finite field extension $L/K$, there is a unique extension of $v$ to a valuation $w$ on $L$; note that in this case $w$ is henselian, too. By [13, Theorem 1.3.1], every complete rank-$1$ valuation is henselian.
@@ 2. $(ii)$
    $u(E)\leqslant u$ for every function field in one variable $E/Kv$.
@@
[…]
@@
Unlike any previous results in the same direction, 1.1 holds for an arbitrary value group $vK$. Independently from our work, a weaker version of the implication from $(ii)$ to $(i)$ was obtained in [21, Theorem 4.1] for the case where $vK\subseteq\mathbb{R}$ (see also 6.3).
We further follow up on the step taken by Claus Scheiderer when extending [15, Cor. 4.12] to [28, Theorem 3] in making the statement more meaningful for real fields by using the definition of the $u$-invariant due to Elman and Lam, which we revisit in Section 4. 1.1 remains exactly the same also in that setting.
In the last part of Section 6, we also include some applications of our results.
An imminent challenge which we do not face in the present article is that of removing the hypothesis that the residue field has characteristic different from $2$. For the local-global principle from [23], this is perfectly possible, as is explained in [2, Theorem 7.1]. The study of quadratic forms over the completion $F^{w}$ of the function field $F$ with respect to a valuation $w$ on $F$, however, poses additional problems in residue characteristic $2$. We leave it to future study whether and how 1.1 can be extended to that case.
@@
The study of quadratic forms over a complete valued field with residue characteristic different from $2$ and value group contained in $\mathbb{R}$ goes back to William Hetherington Durfee [9]. Some of the results were reproven for complete discrete valuations by Tonny Albert Springer [29]. The results from [9] extend to henselian valuations without any condition on the value group. While these statements and techniques are essentially well-known, to the best of our knowledge the literature does not cover them in the desirable generality or under minimal hypotheses. We therefore provide in Section 3 a self-contained and elementary presentation of those statements which are relevant to our topic. We indicate in 3.12 how this compares to the existing literature.
@@ 2. Valuations and torsion-free abelian groups
Let $G$ be a torsion-free abelian group and $H$ a subgroup of $G$ with $r=\mathsf{rrk}(G/H)<\infty$. Let $n\in\mathbb{N}^{+}$. Then $[G:nG]\leqslant n^{r}[H:nH]$. Furthermore, if $G/H$ is finitely generated, then $[G:nG]=n^{r}[H:nH]$.
@@
For any $s\in\mathbb{N}$ and elements $f_{1},\dots,f_{s}\in G$ representing different classes in $G/nG$, denoting by $F$ the subgroup of $G$ generated by $H\cup\{f_{1},\dots,f_{s}\}$, we have that $F/H$ is finitely generated and $f_{1},\dots,f_{s}$ represent distinct classes in $F/nF$, whereby $s\leqslant[F:nF]$. Therefore, to prove that $[G:nG]\leqslant n^{r}[H:nH]$, we may assume without loss of generality that $G/H$ is finitely generated.
@@
Assume that $G/H$ is finitely generated. Then $G/G_{0}$ is finite and $G_{0}\simeq\mathbb{Z}^{r}\times H$, whereby $[G_{0}:nG_{0}]=n^{r}[H:nH]$. Since $G$ is torsion-free, multiplication by $n$ induces an isomorphism $G/G_{0}\to nG/nG_{0}$. In particular $[G:G_{0}]=[nG:nG_{0}]$, whereby $[G_{0}:nG_{0}]=[G:nG]$. Therefore $[G:nG]=n^{r}[H:nH]$. ∎
By an ordered abelian group we mean an abelian group $(G,+,0)$ endowed with a total order relation $\leqslant$ such that, for any $x,y,z\in G$ with $x\leqslant y$, one has $x+z\leqslant y+z$. Note that any ordered abelian group is torsion-free as a group. When $(G,+,0,\leqslant)$ is an ordered abelian group, a convex subgroup is a subgroup $H$ of $G$ such that for all $x\in H$ and $y\in G$ with $0\leqslant y\leqslant x$, we have $y\in H$. One sees that the set of convex subgroups of $G$ is totally ordered by inclusion. The cardinality of the set of proper convex subgroups of $G$ will be called the rank of $G$. By [13, Prop. 3.4.1], the rank of $G$ is bounded from above by $\mathsf{rrk}(G)$.
Given a valuation $v$ on a field $K$, we refer to the rank of the value group $vK$ also as the rank of $v$. Note that an ordered abelian group can be embedded into $\mathbb{R}$ if and only if it is of rank at most $1$. (Consequently, the rank-$1$ valuations on $K$ correspond precisely to the non-trivial absolute values on $K$.)
@@ 3. $(3)$
This follows from [13, Theorem 3.4.3]. ∎
@@
Let $(K,v)$ be a henselian valued field and $K_{0}$ a subfield of $K$. There exists a subfield $K_{1}$ of $K$ containing $K_{0}$ such that $vK_{1}/vK_{0}$ is torsion, $vK/vK_{1}$ is torsion-free, $Kv/K_{1}v$ is algebraic and purely inseparable and $v|_{K_{1}}$ is henselian.
@@
Consider $t\in K$ algebraic over $K_{1}$. Then ${[vK_{1}(t):vK_{1}]\leqslant[K_{1}(t):K_{1}]<\infty}$, so $vK_{1}(t)/vK_{1}$ is torsion, and it follows by the choice of $K_{1}$ that $t\in K_{1}$. Therefore, $K_{1}$ is relatively algebraically closed in $K$.
Consider $t\in K$ transcendental over $K_{1}$. Then $K_{1}(t)/K_{1}$ is a rational function field in one variable, and hence it follows by 2.2 from the choice of $K_{1}$ that the extension $K_{1}(t)v/K_{1}v$ is finite and $vK_{1}(t)/vK_{1}\simeq\mathbb{Z}$. Having this for any $t\in K$, we conclude that the extension $Kv/K_{1}v$ is algebraic and $vK/vK_{1}$ is torsion-free.
@@
Consider $\xi\in Kv$ separable over $K_{1}v$. Since $(K,v)$ is henselian, $\xi$ is the residue of some element $t\in\mathcal{O}_{v}$ algebraic over $K_{1}$. Since $K_{1}$ is algebraically closed in $K$, we obtain that $t\in K_{1}$, whereby $\xi\in K_{1}v$. This shows that $Kv/K_{1}v$ is purely inseparable. ∎

### v2

@@ Abstract.
The $u$-invariant of a field is the largest dimension of an anisotropic quadratic torsion form over the field. In this article we obtain a bound on the $u$-invariant of function fields in one variable over a henselian valued field with arbitrary value group and with residue field of characteristic different from $2$. This generalises a theorem due to Harbater, Hartmann and Krashen and its extension due to Scheiderer. Their result covers the special case where the valuation is discrete. We further give a new proof of a theorem due to Parimala and Suresh bounding by $8$ the $u$-invariant of a function field in one variable over any henselian discretely valued field of characteristic 0 with perfect residue field of characteristic $2$.
Keywords: Quadratic form, henselian valued field, function field in one variable, local-global principle, residue form, real field, torsion form, general $u$-invariant, dyadic valuation
@@ 1. Introduction
The study of numerical field invariants in quadratic form theory was initiated by Irving Kaplansky in 1953. In [18], he introduces three quantities in $\mathbb{N}\cup\{\infty\}$ attached to an arbitrary field of characteristic different from $2$ and suggests to study and compare them. One of them was later called the $u$-invariant, a term which goes back to Hanfried Lenz’ article [24] from 1963.
A field $K$ is called nonreal if 1-1 is a sum of squares in $K$, and otherwise $K$ is called real. For this introduction, let $K$ denote a nonreal field of characteristic different from $2$.
@@
As pointed out in [18, Theorem 3], one always has u(K((t)))=2u(K)u(K(\!(t)\!))=2u(K), where K((t))K(\!(t)\!) denotes the field of Laurent series in $t$ over $K$. Consequently, and since $u(\mathbb{C})=1$, examples of fields of $u$-invariant $2^{r}$ for any given $r\in\mathbb{N}$ are easy to obtain. Incited by this observation, Kaplansky wonders in [18] whether $u(K)$ is always a power of $2$ or $\infty$. This question was answered in the negative in 1989 by Alexander Merkurjev. After giving in [29] a first construction of a field of $u$-invariant 66, he showed in [30] that every positive even integer is the $u$-invariant of some field. Merkurjev’s construction of examples involves direct limits and produces fields which are not finitely generated over any proper subfield. When restricting to fields which are of finite transcendence degree over some familiar base field, no examples are known with a $u$-invariant that is not a power of $2$.
Let $n\in\mathbb{N}$ and let $F/K$ be a field extension of transcendence degree $n$. One may try to determine or at least bound the $u$-invariant of $F$ in terms of $n$ and properties of $K$. Well-known statements of this sort are the following:
  • If $K$ is algebraically closed, then $u(F)\leqslant 2^{n}$.
  • If $K$ is finite, then $u(F)\leqslant 2^{n+1}$.
  • If $K$ is a complete discretely valued field with finite residue field, then $u(F)\leqslant 2^{n+2}$.
The first two bounds were established in 1936 by Chiungtze C. Tsen [39] and independently in 1952 by Serge Lang [22, Theorem 6, Corollary]. The last bound was proven by David Leep [23] in 2013, based crucially on a result on systems of quadratic forms due to Roger Heath-Brown [16]. (When $\mathsf{char}(K)\neq 0$, that bound also follows from [22, Theorem 6 and Theorem 8].)
In each of the three cases above, the upper bound becomes an equality if $F/K$ is finitely generated. This follows by a classical valuation argument; see 4.5 below. Hence, in those situations, we actually have $u(F)=2^{d}$ where $d$ is the cohomological dimension of $F$. This might suggest that $u(F)=2^{n+2}$ would hold as well when $K$ is a nonreal number field and $F/K$ a finitely generated extension of transcendence degree $n$. However, even for $n=1$ this is still an open problem. A more general open question is whether $u(F)$ can be bounded in terms of $u(K)$ and the transcendence degree of $F/K$.
@@
Our main result provides a transfer principle for this common upper bound between a henselian valued field $K$ and its residue field, assuming the latter to be of characteristic different from $2$. In particular, for any odd prime number $p$, applying I below to the field of $p$-adic numbers $\mathbb{Q}_{p}$ with its natural valuation, one retrieves the result from [33] and [17] that $u(F)=8$ for every function field in one variable $F/\mathbb{Q}_{p}$.
Our standard reference for the theory of valued fields is [15]. For a valuation $v$ on $K$, we denote by $Kv$ the residue field and by $vK$ the value group of $v$, and we call $v$ discrete if $vK\simeq\mathbb{Z}$. Given a valuation $v$ on $K$, we call the pair $(K,v)$ a valued field. A valuation $v$ on $K$ is called henselian if, for every finite field extension $L/K$, there is a unique extension of $v$ to a valuation $w$ on $L$; note that in this case $w$ is henselian, too. By [15, Theorem 1.3.1], every complete rank-$1$ valuation is henselian.
@@ 2. $(ii)$
    $u(E)\leqslant u$ for every function field in one variable $E/Kv$ or $vK:2vK]=\infty$.
@@
If $v$ is a discrete valuation, then $[vK:2vK]=2$. Hence I generalises [17, Theorem 4.10], where the equivalence of $(i)$ and $(ii)$ was obtained for the case where $(K,v)$ is a complete discretely valued field with $\mathsf{char}(Kv)\neq 2$. Under the same hypothesis on $(K,v)$, a valuation theoretic local-global principle for isotropy of quadratic forms over function fields in one variable over $K$ was established in [8, Cor. 3.5], from which [17, Theorem 4.10] is easily retrieved. This local-global principle [8, Theorem 3.1] was extended in [28, Cor. 3.19] by the last author of the present article to the situation where the valuation $v$ on $K$ is complete and the value group $vK$ is an arbitrary ordered subgroup of $\mathbb{R}$. It is illustrated in [28, Section 6] how this local-global principle applies to establish upper bounds on the $u$-invariant of function fields in one variable over complete valued fields with value group contained in $\mathbb{R}$. By [3, Theorem 4.4], the local-global principle from [28, Cor. 3.19] extends to the case where $v$ is henselian but not necessarily complete, keeping the assumption that $vK$ is contained in $\mathbb{R}$. Nevertheless, this version of the local-global principle turns out to be applicable to the study of field invariants in the case of function fields in one variable over henselian valued fields without assumption on the value group, by using reduction arguments concerning the rank of the valuation. Such reduction steps have been applied in [3] to the study of the Pythagoras number of function fields, and in [9] to the study of linkage of Pfister forms over function fields. Here we apply them to the study of the $u$-invariant, to prove I.
@@
Unlike any previous results in the same direction, I holds for an arbitrary value group $vK$. Independently from our work, a weaker version of the implication from $(ii)$ to $(i)$ was obtained in [26, Theorem 4.1] for the case where $vK\subseteq\mathbb{R}$ (see also 6.3).
@@
In any case, the applicability of a local-global principle to a problem on field invariants in quadratic form theory depends on the control over the same problem in the local situation, here specifically on the ability to bound the $u$-invariant of a henselian valued field in terms of the $u$-invariant of the residue field. The study of quadratic forms over a complete valued field with residue characteristic different from $2$ and value group contained in $\mathbb{R}$ goes back to William Hetherington Durfee [10]. Some of the results were reproven for complete discrete valuations by Tonny Albert Springer [37]. The results from [10] extend to henselian valuations without any condition on the value group. While these statements and techniques are essentially well-known, to the best of our knowledge the literature does not cover them in the desirable generality or under minimal hypotheses. We therefore include in Section 3 a self-contained and elementary presentation of those statements which are relevant to our topic. We indicate in 3.13 how this compares to the existing literature.
Next to removing any restricting hypothesis on the value group, another aspect giving I a wider scope compared to preceding results on $u$-invariants of function fields is that it applies for the definition of the ‘general $u$-invariant’ introduced by Elman and Lam [12], which is meaningful also for real fields. In this setting, Claus Scheiderer had already extended [17, Cor. 4.12] to [36, Theorem 3], thus covering I in the case of a discrete valuation. In Section 4 we revisit this definition in a way that applies to all fields, also regardless of their characteristic, and we prove some basic bounds for valued fields. In dealing with the $u$-invariant of function fields in one variable, the ‘strong $u$-invariant’ of a field was introduced in [17, Sect. 5] as a short-hand notation. In Section 5, we briefly revisit this notion, basing it on the general $u$-invariant as to fit the study of real and nonreal function fields of any characteristic.
In Section 6, we present the proof of I and then provide some examples of fields where it applies. The obvious challenge left by I is that of removing the restriction on the residue field to be of characteristic different from $2$. A valuation whose residue field has characteristic $2$ is also called dyadic. In Section 7 we do a first step on the problem in the dyadic case. For the local-global principle from [28], this is perfectly possible, as is explained in [2, Theorem 7.1]. The study of quadratic forms over the completion $F^{w}$ of the function field $F$ with respect to a dyadic valuation $w$ on $F$, however, poses additional problems. The $u$-invariant of dyadic complete discretely valued fields was studied in [25]. However, the valuations $w$ appearing in the local-global principle are not all discrete. Using ground work from [14] on quadratic forms over dyadic henselian valued fields, we achieve an alternative proof of the following theorem due to Raman Parimala and Suresh Venapally from [34].
Theorem II (7.11 or [34, Theorem 4.7]).
Let $(K,v)$ be a henselian discretely valued field such that $\mathsf{char}(K)=0$, $\mathsf{char}(Kv)=2$ and $Kv$ is perfect. Let $F/K$ be a function field in one variable. Then $u(F)=8$ if $K$ has a finite field extension of even degree, and otherwise $u(F)=4$.
@@ 2. Valuations and torsion-free abelian groups
Let $G$ be a torsion-free abelian group and $H$ a subgroup of $G$ with $r=\mathsf{rrk}(G/H)<\infty$. Let $n\in\mathbb{N}^{+}$. Then $G:nG]\leqslant n^{r}[H:nH]$. Furthermore, if $G/H$ is finitely generated, then $G:nG]=n^{r}[H:nH]$.
@@
For any $s\in\mathbb{N}$ and elements $f_{1},\dots,f_{s}\in G$ representing different classes in $G/nG$, denoting by $F$ the subgroup of $G$ generated by $H\cup\{f_{1},\dots,f_{s}\}$, we have that $F/H$ is finitely generated and $f_{1},\dots,f_{s}$ represent distinct classes in $F/nF$, whereby $s\leqslant[F:nF]$. Therefore, to prove that $G:nG]\leqslant n^{r}[H:nH]$, we may assume without loss of generality that $G/H$ is finitely generated.
@@
Assume that $G/H$ is finitely generated. Then $G/G_{0}$ is finite and $G_{0}\simeq\mathbb{Z}^{r}\times H$, whereby $G_{0}:nG_{0}]=n^{r}[H:nH]$. Since $G$ is torsion-free, multiplication by $n$ induces an isomorphism $G/G_{0}\to nG/nG_{0}$. In particular $G:G_{0}]=[nG:nG_{0}]$, whereby $G_{0}:nG_{0}]=[G:nG]$. Therefore $G:nG]=n^{r}[H:nH]$. ∎
By an ordered abelian group we mean an abelian group $(G,+,0)$ endowed with a total order relation $\leqslant$ such that, for any $x,y,z\in G$ with $x\leqslant y$, one has $x+z\leqslant y+z$. Note that any ordered abelian group is torsion-free as a group. When $(G,+,0,\leqslant)$ is an ordered abelian group, a convex subgroup is a subgroup $H$ of $G$ such that for all $x\in H$ and $y\in G$ with $0\leqslant y\leqslant x$, we have $y\in H$. One sees that the set of convex subgroups of $G$ is totally ordered by inclusion. The cardinality of the set of proper convex subgroups of $G$ will be called the rank of $G$ and denoted by $\mathsf{rk}(G)$. By [15, Prop. 3.4.1], one has $\mathsf{rk}(G)\leqslant\mathsf{rrk}(G)$.
Given a valuation $v$ on a field $K$, we refer to the rank of the value group $vK$ also as the rank of $v$ and denote it by $\mathsf{rk}(v)$. Note that an ordered abelian group can be embedded into $\mathbb{R}$ if and only if it is of rank at most $1$. (Consequently, the rank-$1$ valuations on $K$ correspond precisely to the non-trivial absolute values on $K$.)
@@ 3. $(3)$
This follows from [15, Theorem 3.4.3]. ∎
@@
Let $(K,v)$ be a henselian valued field and $K_{0}$ a subfield of $K$. There exists a subfield $K_{1}$ of $K$ containing $K_{0}$ such that $vK_{1}/vK_{0}$ is torsion, $Kv/K_{1}v$ is algebraic and purely inseparable, $v|_{K_{1}}$ is henselian and the order of any non-trivial torsion element in $vK/vK_{1}$ is a power of $\mathsf{char}(Kv)$.
@@
Consider $t\in K$ algebraic over $K_{1}$. Then $vK_{1}(t):vK_{1}]\leqslant[K_{1}(t):K_{1}]<\infty$}, so $vK_{1}(t)/vK_{1}$ is torsion, and it follows by the choice of $K_{1}$ that $t\in K_{1}$. Therefore, $K_{1}$ is relatively algebraically closed in $K$.
Consider $t\in K$ transcendental over $K_{1}$. Then $K_{1}(t)/K_{1}$ is a rational function field in one variable, and hence it follows by 2.2 from the choice of $K_{1}$ that the extension $K_{1}(t)v/K_{1}v$ is finite. Having this for any $t\in K$, we conclude that the extension $Kv/K_{1}v$ is algebraic.
@@
Consider $\xi\in Kv$ separable over $K_{1}v$. Since $(K,v)$ is henselian, $\xi$ is the residue of some element $t\in\mathcal{O}_{v}$ algebraic over $K_{1}$. Since $K_{1}$ is algebraically closed in $K$, we obtain that $t\in K_{1}$, whereby $\xi\in K_{1}v$. This shows that $Kv/K_{1}v$ is purely inseparable.
It remains to be shown that the order of any torsion element of $vK/vK_{1}$ is either $1$ or a power of $\mathsf{char}(Kv)$. Hence, for $q\in\mathbb{N}^{+}$ not divisible by $\mathsf{char}(Kv)$ and $t\in{K}^{\times}$ such that $qv(t)\in vK_{1}$, we need to show that $v(t)\in vK_{1}$. Given any such $q$ and $t$, we choose $y\in K_{1}$ such that $v(y)=qv(t)=v(t^{q})$, whereby $yt^{-q}\in{\mathcal{O}}^{\times}_{v}$. Since $Kv/K_{1}v$ is purely inseparable, there exists a positive integer $m$ which is either $1$ or a power of $\mathsf{char}(Kv)$ such that the residue of $(yt^{-q})^{m}$ lies in $K_{1}v$. We fix $u\in{K}^{\times}_{1}$ such that $(yt^{-q})^{m}\in u+\mathfrak{m}_{v}$. Since $(K,v)$ is henselian and $q\neq\mathsf{char}(Kv)$, one has $1+\mathfrak{m}_{v}\subseteq K^{\times q}$. Hence, there exists $z\in{\mathcal{O}}^{\times}_{v}$ with $(yt^{-q})^{m}=uz^{q}$. Hence $(t^{m}z)^{q}=u^{-1}y^{m}\in{K}^{\times}_{1}$. Since $K_{1}$ is relatively algebraically closed in $K$, it follows that $t^{m}z\in{K}^{\times}_{1}$. Since $v(z)=0$, we obtain that $mv(t)=v(t^{m}z)\in vK_{1}$. Since $qv(t)\in vK_{1}$ and $q$ is coprime to $m$, we conclude that $v(t)\in vK_{1}$, as desired. ∎